3. The following data shows the relationship between price and quantity demanded at four different prices for a product:

P = $11, Qd = 16
P = $9, Qd = 24
P = $7, Qd = 32
P = $5, Qd = 40

Using the midpoints formula, what is the price elasticity of demand between: (a) $11 and $9; (b) $9 and $7; (c) $7 and $5?
1
Expert's answer
2020-06-21T19:02:36-0400

A ) "P_1=11,Q_1=16"

"P_2,Q_2=24"

"ed=\frac{16-24}{16+24}\frac2{100}\frac{11-9}{11+9}\frac2{100}"


"=\frac{-8/20}{2/10}"


"=2"

B) "P_1=9,Q_1=24"

"P_2=7,Q_2=32"

"=" "ed=\frac{32-24}{32+24}\frac2{7-9}\frac2{100}"

"=\frac{8/28} {-2/8}"

"=1.14"

B) "P_1=9,Q_1=32"

"P_2=5,Q_2=40"

"=\frac{40-32}{40+32}\frac2{100}\frac{5-7}{5+7}\frac2{100}"


"=\frac{12/36}{-2/6}"

"=1"